# SplitOut: Out-of-the-Box Training-Hijacking Detection in Split Learning via Outlier Detection

### v2

@@ Abstract
Index Terms:
Split learning, training-hijacking, outlier detection.
@@ 1 Introduction
Neural networks often perform better when they are trained on more data using more parameters, which in turn requires more computational resources. However, the required computing resources might be inaccessible in many application scenarios, and it is not always possible to share data freely in fields such as healthcare [2, 35].
Distributed and/or outsourced deep learning frameworks such as split learning [53, 20] and federated learning [5, 29] aim to solve these problems by enabling a group of resource-constrained data holders (clients) to collaboratively train a neural network on their collective data without explicitly sharing private data, and also while offloading a (large) share of computational load to more resourceful servers.
In split learning, a neural network is split into two parts, with the clients computing the first few layers on their private input and sharing the output with the central server, who then computes the rest of the layers. In federated learning, each client computes the full network locally, then their parameter updates are aggregated at a central server, and the aggregated update is shared across the clients. While both of these systems can indeed utilize the collective data set of the clients without raw data sharing, they have been shown to leak information about private data in many different ways [14, 36, 17, 15, 66, 16].
Figure 1: Image reconstruction errors of training-hijacking attack (specifically FSHA [36]) for 4 different datasets. The red line represents the point SplitOut detects the attack, well before the attack converges.
The Problem. What a split learning client learns is fully determined by the server, since the clients’ parameter updates are propagated back from the gradients the server sends. A malicious server can thus launch training-hijacking attacks that have been shown to be effective in inferring clients’ private data [36, 16] and implementing backdoors in the client models [65]. Nevertheless, since training is an iterative process, clients can potentially detect training-hijacking attacks and stop training before the attack converges. Two such detection methods have been proposed in earlier work: an active method that relies on the clients tampering with the data to observe the resulting changes in the system [13], and a passive method that depends only on observing the system [16], motivated by the fact that active methods can be detected and circumvented by the attacker. Both of these methods primarily rely on heuristic approaches that do not fully utilize the clients’ capabilities, which motivates the search for less fragile and simpler-to-use approaches.
@@
Our Contribution. In this work, we show that given modest assumptions on the clients’ compute capabilities, an out-of-the-box outlier detection algorithm can successfully distinguish the gradients sent by a malicious training-hijacking server from those sent by an honest one with almost perfect accuracy, resulting in a simple-yet-effective passive detection method. We empirically demonstrate the effectiveness of the method against the existing training-hijacking attacks [36, 65, 16] and an adaptive attacker that can bypass the active method of [13]. Through the data we collected as part of the experiments, we discover that the differences between the honest training and training-hijacking optimization processes result in training-hijacking gradients having significantly distinguishable neighborhood characteristics compared to honest gradients, which makes the use of the neighborhood-based local outlier factor outlier detection algorithm [6] that can be used out-of-the-box from popular frameworks [39] highly effective. Due to the proactive design of our solution, we think that it can potentially detect some not-yet-known attacks, rather than taking a reactive approach.
We will make the model architectures and training code to reproduce the experiments we report in the paper publicly available.
@@ 2.1 Neural Networks and Data Privacy
Modern machine learning (ML) methods employ the principle of empirical risk minimization [48], where given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ and a space $\mathcal{F}$ of functions parameterized via $\theta$, the goal is to find $\theta^{*}$ which minimizes some loss $\mathcal{L}(f_{\theta},\mathcal{D})$. Many modern systems utilize variants of the iterative stochastic gradient descent (SGD) method, where parameters are updated by sampling a mini-batch $\mathcal{B}\subset\mathcal{D}$ and taking a step in the direction of steepest descent on the loss landscape. In the most basic setting, the update is controlled by the learning rate $\alpha$, such that $\theta^{t+1}\leftarrow\theta^{t}-\alpha\cdot\nabla_{\theta}\mathcal{L}(f_{\theta},\mathcal{B})$.
@@
Neural networks [43] define a certain kind of function space. In its simplest form, each neuron in a neural network layer computes a weighted sum of the previous layer neurons and applies a non-linear activation, such as the sigmoid function. Single-layer neural networks defined in this way are universal approximators [23] provided they are wide enough; however, stacking narrower layers to build deep neural networks (DNNs) leads to more efficient training for the same number of parameters. Practical training of DNNs was made possible by the backpropagation algorithm [44], in which a DNN is treated as a composition of functions and the gradient $\nabla_{\theta}\mathcal{L}$ is decomposed via the chain rule, starting with the final layer and flowing backwards through the network (see e.g. [19] for a definitive treatment of these issues).
@@
Although such defined fully-connected DNNs are highly expressive, this expressivity results in huge search spaces that are difficult to optimize. Instead, the function space is restricted through inductive biases that make assumptions about the structure of the data, such as using convolutional neural networks [31] that are invariant to translations of the features (e.g. edges in images), graph neural networks [28, 52] that are invariant to the order in which the nodes of the graph are processed, or attention-based transformers [49, 11] that are designed to process sequential data such as text. The recent framework of geometric deep learning [7] attempts to formalize these principles.
@@
The effectiveness of these specialized model architectures, along with access to ever more data and compute resources has gradually led to ground-breaking performance in tasks such as language modeling [8] and image generation [42]. This has in turn caused the research community to consider ML systems in a more holistic way, focusing e.g. on their safety [22] and energy use [38], but most importantly for our purposes, data privacy.
@@
Data privacy concerns and DNNs relate in two ways: 1) DNNs often perform better given more data, yet privacy regulations in fields such as healthcare [2, 35] make it difficult to utilize existing data, and 2) neural networks, being the black-boxes they are, can leak training or input data [9, 15]. Such data leakage (see Section 5 for more examples) has led to discussions around whether ML models should be classified as personal data of those whose data was used to train the models [50], and data privacy has been featured as a key concern in very recent regulatory work on artificial intelligence (AI) such as the Blueprint for an AI Bill of Rights published by the US Government [58] and the EU AI Act [51]. Attempts to enable privacy-aware training have resulted in frameworks such as split learning which we focus on next.
@@ 2.2 Split Learning
Split learning (SplitNN) [20, 55, 53] aims to address two challenges: 1) data-sharing in sensitive fields such as healthcare, and 2) training large models in environments with limited compute resources. The main idea is to split a DNN between clients (data holders) and a server. Each client locally computes the first few layers on their private data and sends the output to the server, who then computes the rest of the layers. This way, a group of clients can train a DNN utilizing, but not sharing, their collective data. This framework stands out as being more communication-efficient compared to other collaborative ML systems such as federated learning, especially with a higher number of clients and larger model sizes[46].
In this work, we model the clients as the data holders, and thus SplitNN has two possible settings depending on whether the clients share their labels with the server or not. If they do (Figure 1(a)), the server computes the loss value, initiates backpropagation, and sends the gradients of its first layer back to the client, who then completes the backward pass. If, on the other hand, the clients do not share their labels (Figure 1(b)), the client computes the loss and there are two communication steps instead of one. Finally, clients take turns training with the server using their local data, and each client can obtain the most recent parameters from the last-trained client or through a central weight-sharing server. Interestingly due to this sequential training of clients, having many clients is logically equivalent to having a single client with all the data as the entire model is updated one sample (or mini-batch) at a time.
While a large part of computational work is transferred to the server, this involves a privacy/cost trade-off for the clients. The outputs of earlier layers naturally are closer to the original inputs, and thus “leak” more information. Choosing a split depth is therefore crucial for SplitNN to actually provide data privacy, as has been demonstrated in such attacks [14, 36, 17].
@@ 2.3 Training-Hijacking in Split Learning
A split learning server has full control over the gradients sent to the initial client model and can thus direct it towards an adversarial objective of its choice. We call this attack medium training-hijacking. It has already been exploited to infer the private data of the clients [36, 16] and to implement backdoors in client models [65].
@@ 2.3.1 Feature-Space Alignment
All training-hijacking attacks proposed as of this work [36, 65, 16] share this alignment step, but they differ in how they train $f_{s}$ and the auxiliary models they implement to facilitate the attack.
@@ 2.3.2 Inferring Private Client Data
To infer the private client data $X_{\text{priv}}$ using the feature-space alignment approach as in [36, 16], the shadow model $f_{s}$ is treated as the encoder half of an auto-encoder and the auxiliary model $f_{a}$ acts as the corresponding decoder; i.e. for some input $x$, we have $f_{a}\circ f_{s}(x)\approx x$. Crucially, the auto-encoder is trained on $X_{\text{pub}}$ before the feature-space alignment step in preparation for the attack. Then, provided that $X_{\text{pub}}$ and $X_{\text{priv}}$ are sufficiently similar, the combination $f_{a}\circ f_{c}$ also behaves as an auto-encoder on $X_{\text{priv}}$ and given a client output $z=f_{c}(x)$ for $x\in X_{\text{priv}}$, we have $f_{a}(z)\approx x$. This leads to an inversion attack.
@@ 2.3.3 Implementing Backdoors in the Client Model
A model with a backdoor outputs a different value for the same input when the input is equipped with a backdoor trigger, e.g. an image classifier outputs “dog” when the image is that of a cat but with a small white square added by the adversary [56]. For the attack to work, the model should be able to encode the inputs in a sufficiently different way when they are equipped with the trigger. To launch this attack within the training-hijacking paradigm, [65] assumes that the attacker again has access to some dataset $X_{\text{pub}}$ similar to $X_{\text{priv}}$ but with some of the samples containing the backdoor trigger. The shadow model $f_{s}$ is then combined with two models to form two classifiers: $f_{m}\circ f_{s}$ learns the original classification task, while $f_{a}\circ f_{s}$ learns the classify backdoor samples (i.e. binary classification), with the losses
@@
where $Y_{\text{pub}}\in\{1,...,C\}^{N}$ are the labels for the $C$ classes of the original classification task and $B_{\text{pub}}\in\{0,1\}^{N}$ denote whether an input contains the backdoor trigger or not. Intuitively, $f_{a}$ learns both the classification task (through $L_{m}$) and to encode the backdoor samples (through $L_{a}$). Then finally through the alignment process, $f_{c}$ behaves similarly to $f_{a}$ and also learns to encode the backdoor samples.
@@ 2.4 Detecting Training-Hijacking Attacks
Feature-space alignment is an iterative optimization process, and thus the attacks can be effectively prevented if the clients can detect it before the optimization converges. In the first attempt to do so, [13] observes that the objective in Equation 2 is independent of the label values unlike the classification objective. They propose a method called SplitGuard in which the clients randomly assign some label values and measure how the gradients received from the server differ between the randomized and non-randomized labels. If the server is malicious and is doing feature-space alignment instead of the original task, there should be no significant difference since the loss in Equation 2 does not include the label values. If, on the other hand, the server is honest, the gradients of the loss from the randomized labels should differ significantly from that of the non-randomized labels because success in the original task implies failure in the randomized task and vice versa. The gradients of the randomized labels will direct the model away from the points in the parameter-space that correspond to high performance in the non-randomized task.
To turn this insight into a numerical value, the gradients from the randomized batches are stored in the set $F$ (F for “fake”), and the rest is split into two random subsets $R_{1}$ and $R_{2}$ with $R=R_{1}\cup R_{2}$ (R for “regular”). Then the inequalities
$\theta(F,R)$ $>\theta(R_{1},R_{2})$ (5)
$d(F,R)$ $>d(R_{1},R_{2})$ (6)
should hold where
$d(A,B)$ $=\left|\frac{1}{|A|}\sum_{a\in A}\|a\|-\frac{1}{|B|}\sum_{b\in B}\|b\|\right|$ (7)
$\theta(A,B)$ $=\arccos\left(\bar{A},\bar{B}\right)$ (8)
where $\bar{A},\bar{B}$ denote sums of the elements in the sets $A,B$. These values are then used to compute the SplitGuard score
$S$ $=\frac{\theta(F,R)\cdot d(F,R)-\theta(R_{1},R_{2})\cdot d(R_{1},R_{2})}{d(F,R)+d(R_{1},R_{2})+\varepsilon}$ (9)
$S^{*}$ $=\text{Sigmoid}(\alpha\cdot S)^{\beta}\in(0,1)$ (10)
with $\alpha,\beta$ hyperparameters and $\varepsilon$ a small constant. Based on the inequalities 5, 6, $S^{*}$ is expected to be close to 1 if the server is honest, and significantly lower if the server is malicious.
While the results in [13] indicate that SplitGuard can detect all instances of the feature-space hijacking attack of [36], two weaknesses of the method are that it relies on choosing accurate values for $\alpha$ and $\beta$, and that since it tampers with the training process, it can potentially be detected by the attacker and circumvented.
To overcome the second of these issues, [16] proposes a passive detection mechanism relying on the observation that in a classification task, gradients of the samples with the same label should be more similar than those of the samples with different labels. But since the discriminator loss in Equation 2 does not include the labels, this difference will be less visible if the server is performing feature-space alignment.
Following a paradigm similar to SplitGuard [13], [16] aims to quantify this difference and apply a threshold-based decision rule. For each training step $i$, clients store the pair-wise cosine similarities of the gradients from samples with the same and different labels in $S_{s}(i)$ and $S_{d}(i)$, respectively. The quantification then consists of three components:
@@
    of $S_{s}(i)$ and $S_{d}(i)$, averaged across $n$ training iterations.
@@
where $\hat{G}_{n}=G_{n}^{\lambda_{g}},\hat{E}_{n}=-\log(\lambda_{e}E_{n}+\varepsilon_{e})$, and $\hat{V}_{n}=-\log(\lambda_{v}V_{n}+\varepsilon_{v})$, with the hyperparameters $\lambda_{\text{ds}},\lambda_{g},\lambda_{e},\lambda_{v},\varepsilon_{e},\varepsilon_{v}$, and $\alpha$. Similar to SplitGuard, the abundance of hyperparameters in [16] and the lack of a learnable structure to them make the method fragile, and its generalization to unseen scenarios becomes harder.
In the end, it has been shown that earlier methods can achieve almost-perfect accuracy in detecting training-hijacking attacks given the correct setting, but they give up on the clients’ compute capabilities in favor of heuristic-driven computations. Our main purpose in this work instead is to demonstrate that simpler methods can be utilized instead when clients are assumed to have modest compute power, which is reasonable as data sharing more than efficiency motivates collaborative ML setups in certain scenarios.

### v3

# references.bib

@@ Abstract
Keywords:
Machine learning Data privacy Split learning Training-hijacking
@@ 1 Introduction
Neural networks often perform better when they are trained on more data using more parameters, which in turn requires more computational resources. However, the required computing resources might be inaccessible in many application scenarios, and it is not always possible to share data freely in fields such as healthcare [mercuri_hipaa-potamus_2004].
Distributed and/or outsourced deep learning frameworks such as split learning [vepakomma_split_2018, gupta_distributed_2018] aim to solve these problems by enabling a group of resource-constrained data holders (clients) to collaboratively train a neural network on their collective data without explicitly sharing private data, and also while offloading a (large) share of computational load to more resourceful servers.
In split learning, a neural network is split into two parts, with the clients computing the first few layers on their private input and sharing the output with the central server, who then computes the rest of the layers. While SplitNN can indeed utilize the collective data set of the clients without raw data sharing, SplitNN and other distributed training frameworks have been shown to leak information about private data in different ways [erdogan2021unsplit, pasquini_unleashing_2021, gao2023pcat, fredrikson2015model, zhu2019deep, fuFocusingPinocchioNose2023].
The Problem. What a split learning client learns is fully determined by the server, since the clients’ parameter updates are propagated back from the gradients the server sends. A malicious server can thus launch training-hijacking attacks that have been shown to be effective in inferring clients’ private data [pasquini_unleashing_2021, fuFocusingPinocchioNose2023] and implementing backdoors in the client models [yuHowBackdoorSplit2023]. Nevertheless, since training is an iterative process, clients can potentially detect training-hijacking attacks and stop training before the attack converges. Two such detection methods have been proposed in earlier work: an active method that relies on the clients tampering with the data to observe the resulting changes in the system [splitguard], and a passive method that depends only on observing the system [fuFocusingPinocchioNose2023], motivated by the fact that active methods can be detected and circumvented by the attacker. Both of these methods primarily rely on heuristic approaches that do not fully utilize the clients’ capabilities, which motivates the search for less fragile and simpler-to-use approaches.
@@
Our Solution. We show that given modest assumptions on the clients’ compute capabilities, an out-of-the-box outlier detection algorithm can successfully distinguish the gradients sent by a malicious training-hijacking server from those sent by an honest one with almost perfect accuracy, resulting in a simple-yet-effective passive detection method. We empirically demonstrate the effectiveness of the method against the existing training-hijacking attacks [pasquini_unleashing_2021, yuHowBackdoorSplit2023, fuFocusingPinocchioNose2023] and an adaptive attacker that can bypass the active method of [splitguard]. As a further contribution beyond the method proposed, through the data we have collected during the experiments, we discover that the differences between the honest training and training-hijacking optimization processes result in training-hijacking gradients having significantly distinguishable neighborhood characteristics compared to honest gradients, which explains the effectiveness of the neighborhood-based local outlier factor outlier detection algorithm [breunig2000lof] that can be used out-of-the-box from popular frameworks [scikit-learn]. Due to the proactive design of our solution, we think that it can potentially detect some not-yet-known attacks, rather than taking a reactive approach.
Our work thus results in a unifying framework of training-hijacking attacks in split learning, and leads to new insights on their behavior. We make our code anonymously available at https://github.com/ege-erdogan/splitout.
@@ 2.1 Neural Networks and Data Privacy
Modern machine learning (ML) methods employ the principle of empirical risk minimization [vapnik1991principles], where the aim is to find a function $f$ with parameters $\theta$ that minimize some loss over a dataset. Many modern systems utilize variants of the iterative stochastic gradient descent (SGD) method, where parameters are updated by sampling a batch of data and taking a step in the direction of steepest descent on the loss landscape.
@@
Neural networks [rosenblatt1958perceptron] define a certain kind of function space. In its simplest form, each neuron in a neural network layer computes a weighted sum of the previous layer neurons and applies a non-linear activation, such as the sigmoid function. Single-layer neural networks defined in this way are universal approximators [hornik1991approximation] provided they are wide enough; however, stacking narrower layers to build deep neural networks (DNNs) leads to more efficient training for the same number of parameters. Practical training of DNNs was made possible by the backpropagation algorithm [rumelhart1986learning], in which a DNN is treated as a composition of functions and the gradient $\nabla_{\theta}\mathcal{L}$ is decomposed via the chain rule, starting with the final layer and flowing backwards through the network.
@@
Although such defined fully-connected DNNs are highly expressive, this expressivity results in huge search spaces that are difficult to optimize. Instead, the function space is restricted through inductive biases that make assumptions about the structure of the data, such as using convolutional neural networks [lecun1995convolutional] that are invariant to translations of the features (e.g. edges in images), graph neural networks [kipf2016semi, velivckovic2023everything] that are invariant to the order in which the nodes of the graph are processed, or attention-based transformers [vaswani2017attention, devlin2018bert] that are designed to process sequential data such as text.
@@
The effectiveness of these specialized model architectures, along with access to ever more data and compute resources has gradually led to ground-breaking performance in tasks such as language modeling [brown_language_2020] and image generation [rombach2022high]. This has in turn caused the research community to consider ML systems in a more holistic way, focusing e.g. on their safety [hendrycks2021unsolved] and energy use [patterson2021carbon], but most importantly for our purposes, data privacy.
@@
Data privacy concerns and DNNs relate in two ways: 1) DNNs often perform better given more data, yet privacy regulations in fields such as healthcare [mercuri_hipaa-potamus_2004] make it difficult to utilize existing data, and 2) neural networks, being the black-boxes they are, can leak training or input data [carlini2023extracting, fredrikson2015model]. Data privacy has therefore been featured as a key concern in very recent regulatory work on artificial intelligence (AI) such as the Blueprint for an AI Bill of Rights published by the US Government [White_House_2023] and the EU AI Act [veale2021demystifying]. Attempts to enable privacy-aware training have resulted in frameworks such as split learning which we focus on next.
@@ 2.2 Split Learning
Split learning (SplitNN) [gupta_distributed_2018, vepakomma_no_2018, vepakomma_split_2018] aims to address two challenges: 1) data-sharing in sensitive fields such as healthcare, and 2) training large models in environments with limited compute resources. The main idea is to split a DNN between clients (data holders) and a server. Each client locally computes the first few layers on their private data and sends the output to the server, who then computes the rest of the layers. This way, a group of clients can train a DNN utilizing, but not sharing, their collective data. This framework stands out as being more communication-efficient compared to other collaborative ML systems such as federated learning, especially with a higher number of clients and larger model sizes[singh2019detailed].
In our work, we model the clients as the data holders, and thus SplitNN has two possible settings depending on whether the clients share their labels with the server or not. If they do (Figure 1(a)), the server computes the loss value, initiates backpropagation, and sends the gradients of its first layer back to the client, who then completes the backward pass. If, on the other hand, the clients do not share their labels (Figure 1(b)), the client computes the loss and there are two communication steps instead of one. Finally, clients take turns training with the server using their local data, and each client can obtain the most recent parameters from the last-trained client or through a central weight-sharing server. Interestingly due to this sequential training of clients, having many clients is logically equivalent to having a single client with all the data as the entire model is updated one sample (or mini-batch) at a time.
While a large part of computational work is transferred to the server, this involves a privacy/cost trade-off for the clients. The outputs of earlier layers are closer to the original inputs, and thus “leak" more information. Choosing a split depth is therefore crucial for SplitNN to actually provide data privacy, as has been demonstrated in such attacks [erdogan2021unsplit, pasquini_unleashing_2021, gao2023pcat].
@@ 2.3 Training-Hijacking in Split Learning
A split learning server has full control over the gradients sent to the initial client model and can thus direct it towards an adversarial objective of its choice. We call this attack medium training-hijacking. It has already been exploited to infer the private data of the clients [pasquini_unleashing_2021, fuFocusingPinocchioNose2023] and to implement backdoors in client models [yuHowBackdoorSplit2023].
@@ 2.3.1 Feature-Space Alignment
All training-hijacking attacks proposed as of this work [pasquini_unleashing_2021, yuHowBackdoorSplit2023, fuFocusingPinocchioNose2023] share this alignment step, but they differ in how they train $f_{s}$ and the auxiliary models they implement to facilitate the attack.
@@ 2.3.2 Inferring Private Client Data
To infer the private client data $X_{\text{priv}}$ using the feature-space alignment approach as in [pasquini_unleashing_2021, fuFocusingPinocchioNose2023], the shadow model $f_{s}$ is treated as the encoder half of an auto-encoder and the auxiliary model $f_{a}$ acts as the corresponding decoder; i.e. for some input $x$, we have $f_{a}\circ f_{s}(x)\approx x$. Crucially, the auto-encoder is trained on $X_{\text{pub}}$ before the feature-space alignment step in preparation for the attack. Then, provided that $X_{\text{pub}}$ and $X_{\text{priv}}$ are sufficiently similar, the combination $f_{a}\circ f_{c}$ also behaves as an auto-encoder on $X_{\text{priv}}$ and given a client output $z=f_{c}(x)$ for $x\in X_{\text{priv}}$, we have $f_{a}(z)\approx x$. This leads to an inversion attack.
@@ 2.3.3 Implementing Backdoors in the Client Model
A model with a backdoor outputs a different value for the same input when the input is equipped with a backdoor trigger, e.g. an image classifier outputs “dog" when the image is that of a cat but with a small white square added by the adversary [wang2019neural]. For the attack to work, the model should be able to encode the inputs in a sufficiently different way when they are equipped with the trigger. To launch this attack within the training-hijacking paradigm, [yuHowBackdoorSplit2023] assumes that the attacker again has access to some dataset $X_{\text{pub}}$ similar to $X_{\text{priv}}$ but with some of the samples containing the backdoor trigger. The shadow model $f_{s}$ is then combined with two models to form two classifiers: $f_{m}\circ f_{s}$ learns the original classification task, while $f_{a}\circ f_{s}$ learns the classify backdoor samples (i.e. binary classification), with the losses
@@
where $Y_{\text{pub}}\in\{1,...,C\}^{N}$ are the labels for the $C$ classes of the original classification task and $B_{\text{pub}}\in\{0,1\}^{N}$ denote whether an input contains the backdoor trigger or not. Intuitively, $f_{s}$ learns both the classification task (through $L_{m}$) and to encode the backdoor samples (through $L_{a}$). Then finally through the alignment process, $f_{c}$ behaves similarly to $f_{s}$ and also learns to encode the backdoor samples.
@@ 2.4 Detecting Training-Hijacking Attacks
Feature-space alignment is an iterative optimization process, and thus the attacks can be effectively prevented if the clients can detect it before the optimization converges. In the first attempt to do so, [splitguard] observes that the objective in Equation 2 is independent of the label values unlike the classification objective. They propose a method called SplitGuard in which the clients randomly assign some label values and measure how the gradients received from the server differ between the randomized and non-randomized labels. If the server is malicious and is doing feature-space alignment instead of the original task, there should be no significant difference since the loss in Equation 2 does not include the label values. If, on the other hand, the server is honest, the gradients of the loss from the randomized labels should differ significantly from that of the non-randomized labels because success in the original task implies failure in the randomized task and vice versa. The gradients of the randomized labels will direct the model away from the points in the parameter-space that correspond to high performance in the non-randomized task.
This insight is used to compute the SplitGuard score through the angle and the distances between the two sets of gradients. While the results in [splitguard] indicate that SplitGuard can detect all instances of the feature-space hijacking attack of [pasquini_unleashing_2021], two weaknesses of the method are that it relies on choosing a number of hyperparameters during the computation of the score and the decision threshold, and that since it tampers with the training process, it can potentially be detected by the attacker and circumvented.
To overcome the second of these issues, [fuFocusingPinocchioNose2023] proposes a passive detection mechanism relying on the observation that in a classification task, gradients of the samples with the same label should be more similar than those of the samples with different labels. But since the discriminator loss in Equation 2 does not include the labels, this difference will be less visible if the server is performing feature-space alignment.
Following a paradigm similar to SplitGuard [splitguard], [fuFocusingPinocchioNose2023] aims to quantify this difference and apply a threshold-based decision rule. For each training step $i$, clients store the pair-wise cosine similarities of the gradients from samples with the same and different labels in $S_{s}(i)$ and $S_{d}(i)$, respectively. The quantification then consists of three components:
@@
    of $S_{s}(i)$ and $S_{d}(i)$, averaged across $n$ training iterations with $a,b,c$ the coefficients of the polynomial.
@@
where $\hat{G}_{n}=G_{n}^{\lambda_{g}},\hat{E}_{n}=-\log(\lambda_{e}E_{n}+\varepsilon_{e})$, and $\hat{V}_{n}=-\log(\lambda_{v}V_{n}+\varepsilon_{v})$, with the hyperparameters $\lambda_{\text{ds}},\lambda_{g},\lambda_{e},\lambda_{v},\varepsilon_{e},\varepsilon_{v}$, and $\alpha$. Similar to SplitGuard, the abundance of hyperparameters in [fuFocusingPinocchioNose2023] and the lack of a learnable structure to them make the method fragile, and its generalization to unseen scenarios becomes harder.
In the end, it has been shown that earlier methods can achieve almost-perfect accuracy in detecting training-hijacking attacks given the correct setting, but they give up on the clients’ compute capabilities in favor of heuristic-driven computations. Therefore, the applicability of a less fragile and successful out-of-the-box method is required. Our main purpose in this work is to demonstrate that simpler methods can be utilized instead when clients are assumed to have modest computing power, which is reasonable as data sharing more than efficiency motivates collaborative ML setups in many scenarios.